True or False

To evaluate the integral "\displaystyle{\int \sqrt{x}\ln{(x)}dx}"  by integration by parts,  it is wise to choose "u=\sqrt{x}" and "dv=\ln{(x)}dx"



1
Expert's answer
2021-09-06T15:00:46-0400

False


 it is not wise to choose "u=\sqrt{x}" and "dv=\ln{(x)}dx" .

Since the anti derivative of "\sqrt{x}" , is known but the anti derivative of "lnx" is not. So "u=ln(x)" and "dv=\sqrt{x}dx"

"du=\frac{1}{x}"


"v=\frac{2x^{\frac{3}{2}}}{3}"


This because the derivative of "ln(x)" which is "\frac{1}{x}" , will interact nicely with polynomial terms, whereas the integral is some weird derivative on the order of "xlnx" which looks terrible